OpenStudy (anonymous):

4x^2-4xy+y^2-6=0,whether it's any conic(circle,ellipse,parabola or hyperbola) or not?explain a bit???

OpenStudy (anonymous):

hi sirm

OpenStudy (anonymous):

If x² and y² both have the same coefficients like x² + y² or 3x² + 3y², then it is a circle. If x² and y² both have different coefficients that have the same sign, like 4x² + 9y², or x² + 16y², then it is an ellipse. If x² and y² have different signs, like 25x² - 9y², or 16y² - x², then it is a hyperbola. If the equation has either x² and y², but not both, then it is a parabola. i'm not getting yet(:

OpenStudy (sirm3d):

the general form of a conic section is \[\large Ax^2+Bxy+Cy^2+Dx+Ey+F=0\] the discriminant \[\large B^2-4AC\] can identify the conic section according to this rule: parabola (or denegerate) when \[B^2-4AC=0\] ellipse (or degenerate) when \[B^2-4AC<0\] and hyperbola (or degenerate) when \[B^2-4AC > 0\]

OpenStudy (anonymous):

i have to find diccrimant frist to know whether this is conic or not?

OpenStudy (sirm3d):

yes.

OpenStudy (anonymous):

okay but the equation i wrote i think it's not conicor it is?

OpenStudy (sirm3d):

can you identify A, B and C in the equation above?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A=4,B=-4 and C=1

OpenStudy (sirm3d):

what is the value of the discriminant \[B^2 - 4AC\] is it positive (hyperbola), negative(ellipse), or zero (parabola)?

OpenStudy (anonymous):

it's zero is it?

OpenStudy (anonymous):

parabola that is

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

one question more?\[Y ^{2=4\sqrt{2}}X\]

OpenStudy (anonymous):

how we find discrimant of this type of equation?

OpenStudy (anonymous):

?

OpenStudy (sirm3d):

is this the equation? \[\large y^2=4\sqrt{2}x\]

OpenStudy (anonymous):

no actually the question is\[x^{2}-2xy+y ^{2}-8x-8y=0\]

OpenStudy (anonymous):

''by a rotation of axes,eliminate the xyterm and identify the conic and find its elements?

OpenStudy (anonymous):

i have notes in which this question is solved and there is wriiten like this\[Y ^{2}=4\sqrt{2}\]

OpenStudy (anonymous):

which represent parabola

OpenStudy (sirm3d):

so it is a parabola with line of symmetry the Y-axis

OpenStudy (anonymous):

how we come to know it is parabola?

OpenStudy (anonymous):

without discriminant how u know?

OpenStudy (sirm3d):

the discriminant should be applied when there is an Bxy term, otherwise use the forms of the conic sections parabola: 4a(y-h)=(x-h)^2 or 4a(x-h)=(y-k)^2 ellipse: (x-h)^2/a^2 + (y-k)^2/b^2 = 1 hyperbola: (x-h)^2/a^2 - (y-k)^2/b^2 = 1

OpenStudy (sirm3d):

in standard form (unrotated axis), a parabola has only one squared variable. in an ellipse or hyperbola, both variables are squared. it is an ellipse if the coefficients of x^2 and y^2 have the same sign, hyperbola if they have opposite signs (assuming that both terms are on the same side of the equation of the conic).

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks again